### v1

# Goal-Oriented Semantic Communication for Wireless Image Transmission via Stable Diffusion

Abstract
Efficient image transmission is essential for seamless communication and collaboration within the visually-driven digital landscape. To achieve low latency and high-quality image reconstruction over a bandwidth-constrained noisy wireless channel, we propose a stable diffusion (SD)-based goal-oriented semantic communication (GSC) framework. In this framework, we design a semantic autoencoder that effectively extracts semantic information from images to reduce the transmission data size while ensuring high-quality reconstruction. Recognizing the impact of wireless channel noise on semantic information transmission, we propose an SD-based denoiser for GSC (SD-GSC) conditional on instantaneous channel gain to remove the channel noise from the received noisy semantic information under known channel. For scenarios with unknown channel, we further propose a parallel SD denoiser for GSC (PSD-GSC) to jointly learn the distribution of channel gains and denoise the received semantic information. Experimental results show that SD-GSC outperforms state-of-the-art ADJSCC and Latent-Diff DNSC, with the Peak Signal-to-Noise Ratio (PSNR) improvement by 7 dB and 5 dB, and the Fréchet Inception Distance (FID) reduction by 16 and 20, respectively. Additionally, PSD-GSC archives PSNR improvement of 2 dB and FID reduction of 6 compared to MMSE equalizer-enhanced SD-GSC.
Index Terms:
Goal-oriented semantic communication, Stable diffusion model, and Image transmission.
Figure 1: Stable diffusion-based GSC framework for wireless image transmission
I Introduction
The growing reliance on visual applications in domains such as edge computing and vehicular networks highlights the importance of seamless, reliable, and fast sharing of visual information [1]. To achieve efficient image transmission over bandwidth-limited wireless networks, it is necessary to reduce the transmitted data size while ensuring high-quality image reconstruction [2]. Goal-oriented semantic communication (GSC) has emerged as a promising scheme that focuses on extracting and transmitting the important semantic information relevant to the goal rather than full bits [3]. By selectively transmitting the semantic information, GSC allows for a certain degree of pixel-level error tolerance and evaluates the perception of the reconstructed image, potentially leading to more efficient image transmission [4].
To enhance the efficiency of GSC system, deep learning-based joint source-channel coding (JSCC) schemes have attracted increasing attention. A JSCC-based semantic communication architecture has been developed in [5] for high-resolution image transmission over additive white Gaussian noise (AWGN) and fading channels. Subsequently, a multiple-description JSCC scheme was designed in [6] for bandwidth-agile image transmission. However, these schemes could not understand the semantic content of the transmitted images. To solve this, the work [4] designed a deep reinforcement learning-based semantic encoder and introduced a Generative Adversarial Nets-based semantic decoder with attention mechanism to achieve high-quality image reconstruction. Nonetheless, the aforementioned works assumed that the channel remains consistent during training and inference, which can cause severe performance degradation under practical channel conditions with dynamism. To address this issue, an attention mechanism-based deep JSCC (ADJSCC) scheme was proposed in [7], which can dynamically adjust the SNR during training to adapt to changing channel conditions and remove channel noises. However, adaptive denoising in practical channel conditions with dynamism remains a major challenge.
Generative models, e.g., Denoising Diffusion Probabilistic Models (DDPM) [8], have demonstrated their capability in denoising noisy or imperfect input data across various generation tasks, such as image [9] and video generation [10]. Inspired by this, Latent-Diff DNSC was developed in [2] to characterize noisy channels for efficient image transmission, which demonstrates the effectiveness of DDPM in removing noise from received noisy semantic information. However, this scheme assumes perfect channel estimation for AWGN channels, which poses challenges for accurate denoising in practical channel conditions, especially in the presence of imperfect channel estimation.
To effectively combat the impact of wireless channel noise on semantic information transmission, we propose a stable diffusion (SD)-based GSC framework. The main contributions are summarized in the following:
  • We propose an SD-based GSC framework with a semantic encoder at the transmitter, and a semantic denoiser and a semantic decoder at the receiver. The semantic encoder extracts semantic information from the image; the semantic denoiser denoises the received noisy semantic information; the semantic decoder reconstructs the image from the denoised semantic information.
  • We propose an SD-based semantic denoiser for GSC (SD-GSC) that treats instantaneous channel gain as an input condition of SD to identify and remove noise under known channel. For unknown channel, we propose a parallel SD-based semantic denoiser for GSC (PSD-GSC) that jointly estimates the instantaneous channel gain and denoises the received noisy semantic information.
  • We conduct extensive experiments to examine the Peak SNR (PSNR) and Fréchet Inception Distance (FID). Compared to state-of-the-art ADJSCC and Latent-Diff DNSC, SD-GSC improves the PSNR by 7 dB and 5 dB and reduces the FID by 16 and 20, respectively. Additionally, PSD-GSC improves the PSNR by 2 dB and reduces the FID by 6 compared to the MMSE equalizer-enhanced SD-GSC.
II System Model and Problem Formulation
We consider an image transmission task over Rayleigh fading channels. To achieve efficient image transmission, we introduce a stable diffusion (SD)-based goal-oriented semantic communication (GSC) framework, that consists of a semantic encoder at the transmitter and a semantic denoiser and a semantic decoder at the receiver, as shown in Fig. 1. The semantic encoder extracts the semantic information from the input image, and the semantic decoder then reconstructs the image from the received semantic information. Importantly, the semantic denoiser aims to remove the noise from the received noisy semantic information to enhance the reconstruction capacity of the semantic autoencoder. Specifically, we propose an SD denoiser under known channel, and we further introduce two alternative denoisers for scenarios with unknown channel:
  • MMSE equalizer-enhanced SD (MSD) denoiser: MMSE equalizer is used to estimate the instantaneous channel gain, which is then fed into the SD denoiser,
  • Parallel SD (PSD) denoiser: Two parallel SD modules are used to jointly learn the instantaneous channel gain and denoise the channel noise.
The details of each module are presented one by one below.
At the transmitter, the input data $\bm{x}\in\mathbb{R}^{H\times W\times C}$ is fed into the semantic encoder to extract the semantic information, thereby yielding a latent semantic vector $\bm{z}\in\mathbb{R}^{d}$,
$\bm{z}=\mathcal{E}(\bm{x}),$ (1)
where $\mathcal{E}(\cdot)$ is the operation of semantic encoder.
Subsequently, the semantic information $\bm{z}$ is transmitted through a wireless channel, experiencing transmission impairments such as distortion and noise. The received noisy semantic information is mathematically represented as
$\bm{\bm{z}^{\prime}}=\bm{h}*\bm{z}+\bm{n},$ (2)
where $\bm{h}$ denotes the channel gain between the transmitter and receiver; and $\bm{n}\sim\mathcal{CN}(0,\sigma^{2})$ represents the additive white Gaussian noise (AWGN) with the noise power of $\sigma^{2}$.
At the receiver, the noisy semantic information ${\bm{z}^{\prime}}$ is progressively denoised using the semantic denoiser to obtain a denoised semantic vector as
$\tilde{\bm{z}}=\mathcal{F}_{\epsilon_{\theta}}({\bm{z}^{\prime}}),$ (3)
where $\mathcal{F}_{\epsilon_{\theta}}(\cdot)$ is the denosing operation with parameter $\epsilon_{\theta}$.
The denoised semantic vector $\tilde{\bm{z}}$ is then fed into the semantic decoder to generate the recovered image data $\tilde{\textbf{x}}\in\mathbb{R}^{H\times W\times C}$ as
$\tilde{\bm{x}}=\mathcal{D}(\tilde{\bm{z}}),$ (4)
where $\mathcal{D}(\cdot)$ is the operation of the semantic encoder.
With the aim to ensure the fidelity of image reconstruction, we introduce the mean squared error (MSE) to measure the difference between the transmitted image at the transmitter and the recovered image at the receiver, and formulate the optimization problem as
$\min\frac{1}{N}\sum_{i=1}^{N}\|\bm{x}_{i}-\tilde{\bm{x}}_{i}\|^{2},$ (5)
where $N$ is the size of the image dataset $\mathcal{X}$.
III Architecture of SD-based GSC
In this section, we present our design of semantic autoencoder and semantic denoiser under perfect and imperfect channel estimation, respectively.
III-A Semantic Autoencoder
We design a deep neural network-based semantic autoencoder that can encode input image into a low-dimensional latent semantic space to extract important semantic information. The detailed neural network structure of our designed semantic autoencoder is present in Table I.
To better reconstruct the image, we incorporate MSE and Kullback-Leibler (KL) divergence into the training process of the semantic autoencoder. The MSE facilitates accurate image reconstruction, while the KL divergence serves as a regularization term, guiding the data distribution in the latent space towards approximating a unit Gaussian distribution. The loss function of the semantic autoencoder is defined as the sum of MSE and KL divergence via
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\|\bm{x}_{i}-\tilde{\bm{x}}_{i}\|^{2}+\lambda\mathcal{L}_{\text{KL}},$ (6)
where $\lambda$ is to control the importance of KL divergence with respect to the MSE, and the KL divergence is denoted as
$\mathcal{L}_{\text{KL}}=-\frac{1}{2}\sum_{i=1}^{N}(1+\log(\sigma_{i}^{2})-\mu_{i}^{2}-\sigma_{i}^{2}),$ (7)
where $\mu_{i}$ and $\sigma_{i}$ represent the mean and standard deviation of the latent space, respectively.
III-B SD Denoiser under Known Channel
To handle the wireless channel noise during semantic information transmission, Latent-Diff DNSC used DDPM for image generation. However, the uncontrolled generation process of DDPM may lead to a decrease in image fidelity during reconstruction [9]. Furthermore, the normalization of the received semantic information in Latent-Diff DNSC cannot always guarantee that the normalized data adhere to the standard Gaussian distribution, potentially degrading the reconstruction quality. To mitigate these challenges, we propose an SD Denoiser that uses the received noisy semantic information and instantaneous channel gain as control conditions to effectively remove channel noise, as shown in Fig. 1. Note that SD and DDPM share the same forward process.
III-B1 Forward Process
The forward process (i.e., training process) is executed by iteratively adding Gaussian noise to the initial distribution of semantic information $\bm{z}_{0}\sim p(\bm{z})$ over $T$ time steps to gradually approach an isotropic Gaussian distribution $\bm{z}_{T}\sim\mathcal{N}(0,\mathbf{I})$. In DDPM [11], at time step $t\in[0,T]$, the forward process of $\bm{z}_{t}$ is expressed as the following noisy process:
$\bm{z}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{z}_{0}+\sqrt{(1-\bar{\alpha}_{t})}\epsilon,$ (8)
where $\bm{z}_{0}\sim p(\bm{z})$ is the initial distribution of semantic information, $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\left(1-\alpha_{i}\right)$, and $\beta_{t}$ is the noise scheduling function and is typically modeled as a monotonically increasing linear function of time step $t$, and $\epsilon\sim\mathcal{N}(0,\mathbf{I})$.
From the score-based perspective (i.e., the gradient of the log probability density with respect to the data $\bm{z}_{t}$ at each noise scale $\beta_{t}$), the forward process can be expressed as the following linear stochastic differentiable equation (SDE) [9]:
$\mathrm{d}^{(f)}\bm{z}_{t}=-\frac{\beta_{t}}{2}\bm{z}_{t}\mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}\bm{w},$ (9)
where $\bm{w}$ is the $d$-dimensional Wiener process.
III-B2 Reverse Process
The reverse process (i.e., inference process) is to recover the semantic information $\bm{z}_{0}$ starting from the sample $\bm{z}_{T}$.
DDPM: According to (9), the reverse process of DDPM can be achieved by the corresponding reverse SDE as [8]
$\mathrm{d}^{(r)}\bm{z}_{t}=\big{[}-\frac{\beta_{t}}{2}\bm{z}_{t}-\beta_{t}\nabla_{\bm{z}_{t}}\log p(\bm{z}_{t})\big{]}\mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}\overline{\bm{w}},$ (10)
where $\mathrm{d}t$ is the reverse time step; $\mathrm{d}\overline{\bm{w}}$ is the standard Wiener process in reverse; and $\log p_{t}(\bm{z}_{t})$ is the score function.
Since $\nabla_{\bm{z}_{t}}\log p(\bm{z}_{t})=\nabla_{\bm{\bm{z}}_{t}}\log p\left(\bm{\bm{z}}_{t}\mid\bm{\bm{z}}_{0}\right)$, we can approximate $\nabla_{\bm{z}_{t}}\log p(\bm{z}_{t})\simeq\bm{s}_{\theta}\left(\bm{\bm{z}}_{t},t\right)$ for the reverse process in (10) by solving the following minimization problem during the training in the forward process [12]:
[…]
where the trained score network $\bm{s}_{\theta}\left(\bm{\bm{z}}_{t},t\right)$ can be denoted by using Tweedie’s identity [13] as
$\bm{s}_{\theta}\left(\bm{\bm{z}}_{t},t\right)=\nabla_{\bm{z}_{t}}\log p(\bm{z}_{t})=-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}\left(\bm{\bm{z}}_{t},t\right),$ (12)
where $\epsilon_{\theta}\left(\bm{\bm{z}}_{t},t\right)$ is the learned noise estimator of semantic information with parameter $\theta$ at time step $t$.
TABLE I: Structure of Semantic Autoencoder
Semantic Encoder Semantic Decoder
$\bm{x}\in\mathbb{R}^{H\times W\times C}$ $\bm{z}\in\mathbb{R}^{d}$
$1\times\begin{cases}\text{Conv2d}(3,1,1)\\ \text{ResidualBlock}(C)\\ \text{AttentionBlock}(C)\\ \end{cases}$ Linear: $\mathbb{R}^{h\times w\times c}$
$3\times\begin{cases}\text{Conv2d}(3,2,1)\\ \text{ResidualBlock}(c,c)\\ \text{AttentionBlock}(c)\\ \end{cases}$ $3\times\begin{cases}\text{ConvTranspose2d}(3,2,1,1)\\ \text{ResidualBlock}(c)\\ \text{AttentionBlock}(c)\\ \end{cases}$
BatchNorm, Relu: $\mathbb{R}^{h\times w\times c}$ ConvTranspose2d(3,2,1)
Linear: $\mathbb{R}^{d}$ Sigmoid: $\mathbb{R}^{H\times W\times C}$
SD: Obviously, the above reverse process of DDPM from a random sample $\bm{z}_{0}\sim p(\bm{z})$ cannot ensure the fidelity of image reconstruction. To allow for more controlled and guided image generation, we designed an SD Denoiser conditional on the received noisy semantic information ${\bm{z}^{\prime}}$ and instantaneous channel gain ${\bm{h}}$. Therefore, it is straightforward to modify (10) to derive the reverse process of SD from the posterior distribution
$\mathrm{d}^{(r)}\bm{z}_{t}$ $=[-\frac{\beta_{t}}{2}\bm{z}_{t}-\beta_{t}\nabla_{\bm{z}_{t}}\log p\left(\bm{z}^{\prime}\mid\bm{z}_{t}\right)]\mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}\overline{\bm{w}}$ (13)
$=[-\frac{\beta_{t}}{2}\bm{z}_{t}-\beta_{t}(\nabla_{\bm{z}_{t}}\log p(\bm{z}_{t})$
$\hskip 14.22636pt+\nabla_{\bm{z}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}))]\mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}\overline{\bm{w}}.$
By discretizing the reverse process in (13), we have
$\bm{z}_{t-1}=$ $\frac{1}{\sqrt{\alpha_{t}}}(\bm{z}_{t}+\beta_{t}[\bm{s}_{\theta}\left(\bm{\bm{z}}_{t},t\right)+\nabla_{\bm{z}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h})])$ (14)
$+\sqrt{\beta_{t}}\mathcal{N}(0,\mathbf{I}),$
where $\bm{s}_{\theta}\left(\bm{\bm{z}}_{t},t\right)$ is the trained score network in (12).
To solve the reverse process in (14), the main challenge lies in the posterior distribution $p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h})$. However, the relationship between $\bm{z}^{\prime}$ and $\bm{z}_{0}$ is known but the relationship between $\bm{z}_{t}$ and $\bm{z}^{\prime}$ is unknown. To tackle this issue, we express $p\left(\bm{z}^{\prime}\mid\bm{z}_{t}\right)$ as
$p\left(\bm{z}^{\prime}\mid\bm{z}_{t}\right)=\int p\left(\bm{z}^{\prime}\mid\bm{z}_{0}\right)p\left(\bm{z}_{0}\mid\bm{z}_{t}\right)\mathrm{d}\bm{z}_{0}.$ (15)
where the mean of $p(\bm{z}_{0}\mid\bm{z}_{t})$ can be approximated by a delta function as
$p\left(\bm{z}_{0}\mid\bm{z}_{t}\right)\simeq\delta_{\textbf{E}[\bm{z}_{0}\mid\bm{z}_{t}]}(\bm{z}_{0}),$ (16)
To estimate the $E[\bm{z}_{0}|\bm{z}_{t}]$, we can use the well-trained noise estimator $\epsilon_{\theta}(\bm{z}_{t},t)$ in the forward process (11) to obtain the estimation $E[\bm{z}_{0}|\bm{z}_{t}]=\hat{\bm{z}}_{0}(t)$ as
$\hat{\bm{z}}_{0}(t)=\frac{1}{\sqrt{\alpha_{t}}}(\bm{z}_{t}-\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(\bm{z}_{t},t)).$ (17)
Using (17), the approximation $p(\bm{z}_{0}|\bm{z}_{t})$ leads to the following formula for the gradient of the log-likelihood:
$\nabla_{\bm{z}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h})=-\frac{1-\bar{\alpha}_{t}}{\left(1-\alpha_{t}\right)\left(1-\bar{\alpha}_{t-1}\right)}\nabla_{\bm{z}_{t}}||\bm{z}^{\prime}-\bm{h}*{\bm{z}_{t}}||^{2}.$ (18)
To fasten the convergence of our proposed SD denoiser for GSC (SD-GSC), we use a two-stage training process. We first solely train the semantic autoencoder, then we freeze its parameters and feed the latent semantic vector $\bm{z}$ into the semantic denoiser for training. The training and inference process is detailed in Algorithm 1.
0:  Dataset $\mathcal{X}$, time steps $T$, hyperparameter $\left\{\beta_{1},\cdots,\beta_{T}\right\}$, hyperparameter $\zeta_{t}$, $\lambda$
0:  Reconstructed image $\tilde{\bm{x}}$.
// Training semantic autoencoder
1:  repeat
2:     for each sample $\bm{x}_{i}$ in $\mathcal{X}$ do
3:        Encoding $\bm{x}_{i}$ into a latent vector $\bm{z}_{i}$ use (1).
4:        Decoding $\bm{z}_{i}$ to reconstruct image $\tilde{\bm{x}}$ use (3).
5:     end for
6:     Compute the training loss using (6).
7:     Update the parameter of $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$.
8:  until Converged
9:  Froze parameters of $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$.
// Training semantic denoiser
10:  repeat
11:     Sample a latent representation $\bm{z}_{0}\sim p(\bm{z})$.
12:     $t\sim\textrm{Unifiorm}(\left\{1,2,\cdots,T\right\})$.
13:     $\epsilon\sim\mathcal{N}(0,\mathbf{I})$.
14:     Take gradient descent step on $\nabla_{\theta}\left\|\epsilon_{\theta}\left(\bm{z}_{t},t\right)-\epsilon\right\|_{2}^{2}$.
15:  until Converged
// Denosing process
16:  Sample $\bm{z}_{T}\sim\mathcal{N}(0,I)$
17:  for denosing step $t=T,\ldots,1$ do
18:     $\hat{\epsilon}\leftarrow\epsilon(\bm{z}_{t},t)$.
19:     Estimate $\hat{\bm{z}_{0}}(t)$ using (17).
20:     $g\leftarrow\nabla_{\bm{z}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h})$.
21:     Compute the conditional score $\bm{s}\leftarrow\zeta_{t}g-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\epsilon}$.
22:     Sample $\varphi\sim\mathcal{N}(0,I)$.
23:     Compute $\bm{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{z}_{t}+\beta_{t}\bm{s})+\sqrt{\beta_{t}}\varphi$
24:  end for
Algorithm 1 SD-GSC under known channel
III-C PSD Denoiser under Unknown Channel
Notably, Algorithm 1 is only applicable when the instantaneous channel gain $\bm{h}$ is known, and hence cannot be directly used for the scenarios (e.g., Massive MIMO communications systems) with imperfect estimation of $\bm{h}$. To solve this issue, we propose a parallel SD-GSC (PSD-GSC) to jointly estimate the channel gain and remove the noise.
III-C1 Forward Process
Since $\bm{z}$ and $\bm{h}$ are independent, the posterior probability is given by
$p(\bm{z},\bm{h}|\bm{\bm{z}^{\prime}})=p(\bm{\bm{z}^{\prime}}|\bm{z},\bm{h})p(\bm{z})p(\bm{h}).$ (19)
Therefore, we can train two separate forward processes for $\bm{z}$ and $\bm{h}$, respectively. Similar to (12) and (17), the score network of $\bm{h}$ and $E[\bm{h}_{0}|\bm{h}_{t}]=\hat{\bm{h}}_{0}(t)$ can be estimated using
$\bm{s}_{\vartheta}\left(\bm{\bm{h}}_{t},t\right)=\nabla_{\bm{h}_{t}}\log p(\bm{h}_{t})=-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\vartheta}(\bm{h}_{t},t),$ (20)
and
$\hat{\bm{h}}_{0}(t)=\frac{1}{\sqrt{\alpha_{t}}}(\bm{h}_{t}-\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\vartheta}(\bm{z}_{t},t)),$ (21)
where $\epsilon_{\vartheta}\left(\bm{\bm{h}}_{t},t\right)$ is the learned noise estimator of channel gain with parameter $\theta$ at time step $t$.
III-C2 Reverse Process
Similar to the reverse process in (13), the two reverse processes of PSD-GSC can be expressed as
$\mathrm{d}^{(r)}\bm{z}_{t}=$ $\big{[}-\frac{\beta_{t}}{2}\bm{z}_{t}-\beta_{t}(\nabla_{\bm{z}_{t}}\log p(\bm{z}_{t})$ (22)
$+\nabla_{{\bm{z}}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}_{t}))\big{]}\mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}\overline{\bm{w}},$
and
$\mathrm{d}^{(r)}\bm{h}_{t}=$ $\big{[}-\frac{\beta_{t}}{2}\bm{h}_{t}-\beta_{t}(\nabla_{\bm{h}_{t}}\log p(\bm{h}_{t})$ (23)
$+\nabla_{{\bm{h}}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}_{t}))\big{]}\mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}\overline{\bm{w}}.$
Correspondingly, the above two reverse processes can be discretized in a similar way as (14) and expressed as
$\bm{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{z}_{t}+$ $\beta_{t}[\nabla_{\bm{z}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}_{t})$ (24)
$+\bm{s}_{\theta}\left(\bm{\bm{z}}_{t},t\right)])+\sqrt{\beta_{t}}\mathcal{N}(0,\mathbf{I}),$
and
$\bm{h}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{h}_{t}+$ $\beta_{t}[\nabla_{\bm{h}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}_{t})$ (25)
$+\bm{s}_{\vartheta}\left(\bm{\bm{h}}_{t},t\right)])+\sqrt{\beta_{t}}\mathcal{N}(0,\mathbf{I}).$
Due to the sparse structure of wireless channel, we use $\ell_{1}$ regularization to sparse the channel gain by augmenting the diffusion prior thereby better stabilize the reconstruction. The minimization strategy for the channel gain then becomes
$\bm{h}^{\prime}_{t-1}=\bm{h}_{t-1}-\alpha(||\bm{\bm{z}^{\prime}}-\bm{h}_{t-1}*\bm{z}_{t-1}||_{2}+\phi||\bm{h}_{t-1}||),$ (26)
where $\phi$ is the regularization strength.
The detailed process of our proposed PSD denoiser for GSC (PSD-GSC) under unknown channel is presented in Algorithm 2.
0:  Training dataset $\mathcal{X}$, learning rate $\gamma$, time steps $T$, hyperparameter $\left\{\beta_{1},\cdots,\beta_{T}\right\}$, $\zeta_{t}^{\theta}$, $\zeta_{t}^{\vartheta}$ and $\phi$
0:  Reconstructed image $\tilde{\bm{x}}$.
1:  Train semantic autoencoder $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$ using steps 1-8 in Algorithm 1, then freeze the parameters.
// Training semantic denoiser
2:  Train denoising estimator $\epsilon_{\theta}$ and $\epsilon_{\vartheta}$ for $\bm{z}$ and $\bm{h}$ following steps 10-15 in Algorithm 1, respectively.
// Denosing process
3:  Sample $\bm{z}_{T},\bm{h}_{T}\sim\mathcal{N}(0,I)$.
4:  for denosing step $t=T,\ldots,1$ do
5:     $\hat{\epsilon}_{\theta}\leftarrow\epsilon_{\theta}(\bm{z}_{t},t)$, $\hat{\epsilon}_{\vartheta}\leftarrow\epsilon_{\vartheta}(\bm{z}_{t},t)$.
6:     Estimate $\hat{\bm{z}}_{t}$ and $\hat{\bm{h}}_{t}$ using (17) and (21).
7:     $g_{\theta}\leftarrow\nabla_{\bm{z}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}_{t})$, $g_{\vartheta}\leftarrow\nabla_{\bm{h}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}_{t})$.
8:     Compute the conditional score
$\small\bm{s}_{\theta}\leftarrow\zeta_{t}^{\theta}g_{\theta}-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\epsilon}_{\theta}$ and $\bm{s}_{\vartheta}\leftarrow\zeta_{t}^{\vartheta}g_{\vartheta}-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\epsilon}_{\vartheta}$.
9:     Sample $\varphi_{\theta},\varphi_{\vartheta}\sim\mathcal{N}(0,I)$.
10:     Compute $\bm{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{z}_{t}+\beta_{t}\bm{s}_{\theta})+\sqrt{\beta_{t}}\varphi_{\theta}$ and $\bm{h}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{h}_{t}+\beta_{t}\bm{s}_{\vartheta})+\sqrt{\beta_{t}}\varphi_{\vartheta}$.
11:     ${\bm{h}_{t-1}\leftarrow\bm{h}_{t-1}-\alpha(||\bm{\bm{z}^{\prime}}-\bm{h}_{t-1}*\bm{z}_{t-1}||_{2}+\phi||\bm{h}_{t-1}||)}$.
12:  end for
Algorithm 2 PSD-GSC under unknown channel
IV Performance Evaluation
(a) PSNR
(b) FID
Figure 2: PSNR and FID under known channel
We implement the image transmission task under a Rayleigh fading channel, using the FFHQ dataset [14], which consists of 70,000 high-quality human face images at a resolution of 256×256 pixels, that is created as a benchmark for generative tasks. In the implementation of the semantic autoencoder, the encoder downsamples the feature dimensions by half, starting from the 2nd block, with channel dimensions specified as $[64,128,256,512]$, and then the features are embedded into a latent representation of dimension $d=4096$. Conversely, the decoder upsamples the feature dimensions by double for the first 3 blocks, with channel dimensions set to be $[512,256,128,64]$. For the semantic denoiser, we set the time step $T=1000$ and use a linear variance scheme to determine the hyperparameters of $\left\{\beta_{1},\beta_{2},\cdots,\beta_{T}\right\}$.
Baselines: To demonstrate the effectiveness of our proposed frame, we perform a comparative analysis against the state-of-the-art baselines: ADJSCC [7] and Latent-Diff DNSC [2]. We also compare with MSD denoiser-based GSC (MSD-GSC) as a benchmark for channel gain estimation to highlight the effectiveness of our PSD-GSC in learning channel gains.
Metrics: We evaluate two performances: peak signal-to-noise ratio (PSNR) to evaluate image reconstruction quality, with higher values indicating better reconstruction quality; and Frechet inception distance (FID) to assess the similarity between real and reconstructed image distributions, with lower values indicating better reconstruction quality.
IV-A Known Channel
Fig. 2 presents the PSNR and the FID performance across various SNRs under known channel. As expected, both PSNR and FID performances improve as the SNR increases, owing to the fact that better wireless channel conditions facilitate higher-quality image reconstruction. Interestingly, the Latent-Diff DNSC scheme outperforms the ADJSCC scheme in both PSNR and FID at all SNRs, due to diffusion model can efficiently capture the wireless channel characteristics and remove the noise from the received noisy semantic information, resulting in a more accurate reconstruction. Importantly, our proposed SD-GSC outperforms two baseline schemes in both PSNR and FID metrics. Specifically, compared to Latent-Diff DNSC and ADJSCC, the PSNR of our proposed SD-GSC achieves approximately 5 dB and 7dB improvement, and the FID of our proposed SD-GSC achieves around 16 and 30 reduction, respectively. This is because integrating the instantaneous channel gain will enable better characterization of channel gain and mitigation of distortion due to channel noise, resulting in a more controlled and guided image generation than the uncontrolled generation process of DDPM used in the Latent-Diff DNSC scheme.
IV-B Unkown Channel
To validate the effectiveness of our proposed PSD-GSC for channel estimation as well as efficient image transmission, we also compare our proposed PSD-GSC with MSD-GSC and SD-GSC. Fig. 3 plots the PSNR and the FID for various SNRs under unknown channel. As expected, our proposed SD-GSC with known $h$ outperforms PSD-GSC and MSD-GSC with unknown $h$ in terms of both PSNR and FID, due to the errors in unknown channel estimation. Specifically, our proposed PSD-GSC and MSD-GSC experience degradation in PSNR by approximately 1 dB and 3 dB, and degradation in FID by 2 and 8, respectively. This is because channel estimation errors result in inaccurate channel gain inputs of the diffusion model, which decreases its denoising and high-quality image reconstruction capabilities. Importantly, PSD-GSC outperforms MSD-GSC both in PSNR and FID. This is due to the probabilistic modeling-based diffusion model used in PSD-GSC can well estimate nonlinear channels and capture the complex properties of wireless channels, for better image reconstruction.
(a) PSNR
(b) FID
Figure 3: PSNR and FID under unknown channel
V Conclusion
In this paper, we developed a stable diffusion (SD)-based goal-oriented semantic communication framework (SD-GSC) for efficient image transmission over Rayleigh fading channels. The main goal is to reduce the bandwidth requirement while ensuring high-quality image reconstruction at the receiver. To this end, we first designed a semantic autoencoder to extract semantic information to reduce the transmitted data size while accurately reconstructing the image from the received semantic information. To combat the negative impact of wireless channel noise on the transmitted semantic information, we designed an SD-based semantic denoiser to better characterize and mitigate the channel noise, leading to significant improvements in both PSNR and FID metrics compared to the state-of-the-art baselines under known channel. For unknown channel scenarios, we further designed a parallel SD-GSC (PSD-GSC) to jointly learn the channel gain and denoise the received semantic information, showcasing the capability of our proposed PSD-GSC in channel estimation and efficient image transmission.